Lockeran Co. has the following projected sales, costs, net investment and free cash flows in millions. The anticipated growth rate in free cash flows after year 6 is 3% per year forever. there are 3 million shares outstanding and investors require a return of 8% on the company's stock.

1.Using the constant growth model to find the terminal value, calculate the price of the company's stock.

2. Suppose instead that you estimate the terminal value using a PE multiple of 0.75, calculate the price of the company's stock.

1

2

3

4

5

6

Sales

176.00

200.64

224.72

247.19

266.97

282.99

Costs

82.5

94.05

105.34

115.87

125.15

132.66

Taxes

17.33

19.75

22.12

24.33

26.28

27.85

Net Income (OCF)

76.17

86.84

97.26

106.99

115.54

122.48

Net Investment

45

51.3

57.46

63.21

68.27

72.37

FCF

31.17

35.54

39.80

43.78

47.27

50.11
Answer #1
1 2 3 4 5 6
Sales 176 200.64 224.72 247.19 266.97 282.99
Costs 82.5 94.05 105.34 115.87 125.15 132.66
Taxes 17.33 19.75 22.12 24.33 26.28 27.85
Net Income (OCF) 76.17 86.84 97.26 106.99 115.54 122.48
Net Investment 45 51.3 57.46 63.21 68.27 72.37
FCF 31.17 35.54 39.8 43.78 47.27 50.11
PV of Formula =31.17/(1+8%)^1 =35.54/(1+8%)^2 =39.8/(1+8%)^3 =43.78/(1+8%)^4 =47.27/(1+8%)^5 =50.11/(1+8%)^6
PV of FCF 28.86 30.47 31.59 32.18 32.17 31.58
PV of all FCF 186.85
Terminal value using model
Terminal value = (FCF*(1+growth rate))/(cost of capital - growth rate)
Terminal value =(50.11*(1+3%))/(8%-3%)
Terminal value = 1032.27
PV of Terminal value = 1032.27/(1+8%)^6
PV of Terminal value = 650.51
Total value = PV of all FCF + PV of Terminal value
Total value = 186.85+650.51
Total value = 837.36
Terminal value using P/E Multiple
Terminal value = (P/E )* (Year 6 Net income)
Terminal value = 0.75*122.48
Terminal value = 91.86
PV of Terminal value =91.86/(1+8%)^6
PV of Terminal value = 57.89
Total value = PV of all FCF + PV of Terminal value
Total value = 186.85+57.89
Total value = 244.74